OpenStudy (anonymous):

a piggy bank contains $2.25.the piggy bank contains only quarters and dimes.There are twice as many dimes as quarters in the bank.How many dimes are in the bank?

OpenStudy (anonymous):

A quarter is one fourth, 1⁄4, or 25% D=0.10 X(0.10) + Y(0.25) = 2.25 X = 2Y Now, replace the X in the longer equation with its known Y equivalent. 2Y(0.10) + Y(0.25) = 2.25 Multiply out the brackets 0.2Y + 0.25Y = 2.25 Add the Y values together 0.45Y = 2.25 Divide both sides by 0.45 Y = 5 So, now we can replace the Y in the longer equation X(0.10) + 5(0.25) = 2.25 and resolve that X(0.10) + 1.25 = 2.25 X(0.10) = 2.25 - 1.25 X(0.10) = 1.00 X = 10 So, there are ten dimes and five quarters